Pete weighs 20 kg more than Beth while Henry weighs 30 kg less than twice as much as Beth. if Keith and Henry weigh the same, how much does Beth weigh? (in kg)
Answer by jokaehler(26)

If Pete (let's call him "x") weighs 20kg more than Beth ("y"), and Henry ("z") weighs 30kg less than twice as much as Beth how much does Beth weigh if Pete and Henry weigh the same?
From this, we can come up with...

x = y + 20
(Because Pete [x] weighs 20kg more than Beth [y])

z = 2y - 30
(Because Henry [z] weighs 30kg less than twice as much as Beth [y])


If Pete and Henry weigh the same, then x = z; therefore...

y + 20 = 2y - 30
y + 50 = 2y
50 = y


Since "y" is Beth, Beth weighs 50kg!